Set-up the integral for the area A(R) of the region R bounded by the circles y


2 + x


2 = 1, and y


2 + x


2 = 4,


below the x-axis and to the left by the y-axis using the vertical strip.

1
Expert's answer
2021-11-01T19:30:47-0400
"A=\displaystyle\int_{-2}^{-1}(0-(-\sqrt{4-x^2}))dx"

"+\displaystyle\int_{-1}^{0}(-\sqrt{1-x^2}-(-\sqrt{4-x^2}))dx"

"=\displaystyle\int_{-2}^{0}\sqrt{4-x^2}dx-\displaystyle\int_{-1}^{0}\sqrt{1-x^2}dx"

"=\dfrac{1}{4}\pi(2)^2-\dfrac{a}{4}\pi(1)^2=\dfrac{3\pi}{4} ({units}^2)"